direcaly douunldod on a thin, wedg e-shpe Fim with no Of refraction 33. An obsrveotny dun ough ths op plate Sees bright and ik frisdeu thth end shese to plates touch? wy? ( Fiva the im ick 04 the 4 bright fringe? (c)Find the difRhence in hickness ueen the adjacent dk agreye
Answer #1

4) Wedge shape film of index of refraction nw = 1.33

Light refecting from top surface undergoes phase change \pi , equivalent path difference 0.5\lambda

Light reflecting from bottom surface undergoes zero phase and path difference

Total path difference between top and bottom surface = 0.5\lambda

Thus condition for maxima or bright fringe

Path difference inside film 2L = (n+(1/2)\lambdan) where, n= 0,1,2,3..... and \lambda n= wavelength inside film

  2L = (n + \frac{1}{2})\frac{\lambda }{n_{w}}..................................................1)

where nw = 1.33 , refractive index of medium, L = thickness of the film

Condition for minima or dark fringe

2L = (n) \frac{\lambda }{n_{w}} ................................................2)

a) when L = 0

Total Path difference of waves reflecting from top and bottom surface

= 0.5 \lambda

Thus waves are out of phase with each other. which produces dark fringe

b) \lambda = 630 nm

At 4th bright fringe n = 4

Substituting value of \lambda and n in equation 1

2L = (4 + \frac{1}{2})\frac{630\times 10^{-9} }{1.33}

  L = 1.065\mu m

c) Difference in thickness for adjacent bright and dark fringe

for n =1 ,

  2(L_{1d}-L_{1b})= \frac{\lambda }{n_{w}}-\frac{\lambda }{2n_{w}}

L_{1d}-L_{1b}= \frac{\lambda }{4n_{w}}

L_{1d}-L_{1b}= \frac{630\times 10^{-9} }{4\times 1.33}

L_{1d}-L_{1b}= 0.118\mu m